# The Primordial Inflation Explorer (PIXIE): Mission design and science goals

### v1

@@ 1 Introduction
Observations of cosmological radiation backgrounds have played a key role in our understanding of the universe. Measurements of the CMB frequency spectrum [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] are consistent with a single blackbody over a range of 3 decades in frequency and 4 decades in intensity (Figure 1). Data from the COBE/FIRAS spectrometer determine the monopole temperature $T_{0}=2.72548\pm 0.00057$ K [26]. The FIRAS limits on deviations from a blackbody spectrum (spectral distortions) at levels $\Delta I/I<5\times 10^{-5}$ [25] support the thermal hot big bang model while ruling out alternatives such as steady state models [27]. Measurements of CMB anisotropies in temperature and polarization have provided insight into the contents of the universe and their evolution from primordial density perturbations to matter clustering, reionization, and the growth of large scale structure, consistent with a single 7-parameter cosmological model ($\Lambda$CDM, [28, 29] ).
@@
The Primordial Inflation Explorer (PIXIE) is a proposed space mission to measure the spectrum and polarization of the cosmic microwave background and astrophysical foregrounds on angular scales of 1 and larger [36, 37, 38]. Operating from the second Sun-Earth Lagrange point (L2), it would survey the entire sky in intensity and linear polarization at integrated sensitivity of a few Jy sr-1 in 300 spectral channels from 28 GHz to 6 THz (Figure 2). In this paper, we describe the PIXIE instrument and highlight the anticipated scientific results from its baseline 2-year mission.
@@ 2 PIXIE Mission
Figure 3: Left: Exploded view of the PIXIE observatory. Nested, passively cooled shells surround a polarizing Fourier transform spectrometer, which is actively cooled to $2.725$ K to match CMB emission. Right: PIXIE’s optical design measures both intensity and polarization from microwave to THz frequencies. As the central mirror pair moves, the detectors measure a fringe pattern proportional to the Fourier transform of the difference between one linear polarization from beam A and the orthogonal polarization from beam B.
The PIXIE instrument (Figure 3) is a polarizing Fourier transform spectrometer (FTS) with two input ports and two output ports. The input ports are fed by two off-axis telescopes that produce twin beams co-pointed along the spacecraft spin axis. Each telescope has a primary mirror $550$ mm in diameter. Folding flats and secondary mirrors route the beams to the FTS while retaining orthogonal polarization alignment between the two beams. Within the FTS a set of five transfer mirror pairs, each imaging the previous mirror to the following one, shuttles the radiation through a series of polarizing wire grids. Polarizer A transmits one polarization and reflects the other, separating each beam into orthogonal polarization states Polarizer B, oriented $45^{\circ}$ relative to Polarizer A, mixes the polarization states. A Mirror Transport Mechanism (MTM) moves the central mirror pair ±$\pm 4$ mm to produce an optical phase delay of up to ±$\pm 15$ mm. The phase-delayed beams re-combine (interfere) at Polarizer C. Polarizer D (aligned to Polarizer A) splits the beams again and routes them to two multi-moded concentrators. Each concentrator has a square aperture to preserve linear polarization [39, 40] and contains a pair of orthogonal polarization-sensitive bolometers mounted back-to-back. The square aperture is rotated by 45 with respect to the plane of polarization from Polarizer D to enforce symmetry between orthogonal linear polarizations [40]. The four bolometers are cooled to 0.1 K [41] to achieve background-limited sensitivity. The optical spectrum is the Fourier conjugate of the detector’s time-ordered data across the mirror sweep. The mirror stroke length sets the optical frequency resolution while the detector sampling rate determines the highest optical frequency.
@@
The PIXIE instrument and mission have evolved since the initial 2011 concept [36]. The 2011 version assumed observations from a polar sun-synchronous low-Earth orbit, while the current version will observe from the second Sun-Earth Lagrange point. The optical path through the FTS has been shortened, replacing the previous corner reflector MTM with a pair of powered mirrors. The sensitivity curves now include the frequency-dependent effects of the reflective backshort in the detector integrating cavity.
The baseline 2-year mission assumes that 45% of observing time is spent with the calibrator deployed (sensitive to spectral distortions and polarization) and 45% with the calibrator stowed (sensitive to polarization only). The remaining 10% conservatively accounts for time not used for science observations (MTM turnaround at the end of each stroke, thermal settling after setpoint changes, etc,). The calibrator deployment schedule can be altered during the mission to provide more or less integration time in each mode. Mission operations further assume observations at two different MTM stroke lengths, providing additional sensitivity to low-frequency foregrounds.
@@ 2.1 Sensitivity
The PIXIE channel width is determined by the need to separate cosmological signals from competing foregrounds (§$\S$2.2). Channel widths $\Delta\nu\sim 20$ GHz providing spectral resolution $\nu/\Delta\nu\sim 15$ near the CMB peak at 270 GHz do not provide sufficient resolution to fit multiple foreground components (synchrotron, free-free, and anomalous microwave emission) at frequencies below 70 GH, while channel widths $\Delta\nu<10$ GHz appropriate for low-frequency foregrounds degrade the sensitivity to CMB signals at higher frequencies. As a compromise, PIXIE spends 70% of the observing time using broad channel widths for continuum CMB signals, and 30% with narrow widths for low-frequency foregrounds. It is convenient to set all channel widths commensurate with the CO J=1-0 line at 115.3 GHz within the Galaxy, $\Delta\nu=\nu_{\rm CO}/M$ so that every $M^{\rm th}$ channel is centered on a Galactic CO line. For measurements of spectral distortions, PIXIE will use $M=6$ ($\Delta\nu=19.2$ GHz) and $M=12$ ($\Delta\nu=9.6$ GHz). The free-free and anomalous microwave foreground emission are not polarized at relevant levels, allowing use of wider channel widths $M=3$ ($\Delta\nu=38.4$ GHz) and $M=6$ ($\Delta\nu=19.2$ GHz) for polarization measurements. Appendix C lists the resulting channel sensitivities.
Figure 6: Solid lines show the spectra of the principal foregrounds at the median amplitude for Galactic latitude $|b|>20^{\circ}$. Dotted and dashed lines show the cosmological signals, with negative amplitudes dashed. The black/red curve shows the PIXIE noise in each spectral channel for the baseline 2-year mission, averaged over the cleanest 70% of the sky. PIXIE’s sensitivity to cosmological signals is limited by competing emission from astrophysical foregrounds (see text).
@@ 2.2 Foreground Subtraction
Table 1 shows the signal to noise ratio for the cosmological signals estimated using a Fisher analysis of the noise levels integrated over the cleanest 70% of the sky. We adopt fiducial values $y=1.77\times 10^{-6}$ and $kT_{e}=1.245$ keV predicted for the integrated signal from Compton scattering from groups and clusters of galaxies [51] and $\mu=2\times 10^{-8}$ for the dissipation of primordial density perturbations from a power-law distribution consistent with the scale-invariant spectrum seen in primary CMB anisotropies [52]. Results are shown for the baseline PIXIE mission (two years of observations with 45% of observations devoted to spectral distortions, yielding 7.6 months of integration for monopole distortions within the cleanest 70% of the sky) as well as an extended mission with a total of 7 years integration time.
@@
In the absence of foreground emission, the baseline PIXIE mission would detect the cosmological signals at high confidence: 3101 standard deviations for $y$, 80 standard deviations for $kT_{e}$, and 2.9 standard deviations for $\mu$. The need to identify and subtract foreground emission degrades these values. A minimal fit using just the PIXIE mission data with no foreground priors provides a robust detection of the predicted Compton signal (18 standard deviations for $y$ and 2 standard deviations for $kT_{e}$) but falls short of a detection of the chemical potential predicted in the standard $\Lambda$CDM model from Silk dampling of a power-law distribution of primordial density perturbations (§$\S$3.2). We investigate the ability to improve these limits by repeating the analysis while placing priors on selected foreground parameters. PIXIE data at frequencies above $\sim$600 GHz place tight constraints on the amplitude and spectra of the thermal dust and CIB foregrounds and are unlikely to be superseded by other data. The sensitivity at lower frequencies is limited by the fixed channel width imposed by Fourier transform spectroscopy (§$\S$2.1) and can be improved by including information from measurements at lower frequencies. We repeat the Fisher analysis placing Gaussian priors of 10% or 1% on the synchrotron and free-free parameters. Placing 10% priors on the low-frequency foregrounds provides a five-fold improvement in the sensitivity to cosmological signals, but still falls short of a detection for the minimal $\mu$ distortion from a $\Lambda$CDM cosmology. Even the combination of tighter (1%) priors over a 7-year extended mission produces uncertainty larger than the predicted chemical potential distortion. Obtaining additional sensitivity to low-frequency foregrounds would require additional instrumentation to reduce photon noise at frequencies below $\sim$150 GHz, extend measurements to frequencies below 30 GHz, or both. §$\S$4 discusses possible enhancements for such a “Super-PIXIE” design.
@@
Foreground subtraction in polarization is less complex, as only the synchrotron and thermal dust foregrounds are expected to have significant polarization. Figure 7 shows PIXIE’s polarization capability using a Needlet Internal Linear Combination (NILC) component separation [53, 54, 55, 56]). We generate a noiseless full-sky foreground model using the Python Sky Model d1s1 variant [57] and add cosmological signals from E-modes, lensing B-modes, and primordial B-modes at amplitude $r=5\times 10^{-3}$. All signals are convolved with a 1.6 full width at half maximum Gaussian beam before adding noise within each PIXIE frequency channel. We extend the PIXIE frequency coverage using the Planck polarization maps, then evaluate the resulting sky maps on the cleanest 70% of the sky, using a Fisher analysis to estimate the uncertainty $\sigma(r)$ and a likelihood analysis to determine the 95% confidence upper limit $r_{95}$ [58]. As with spectral distortions, the foreground subtraction increases the effective noise with noise amplitude (after component separation) worse by a factor 1.8 in amplitude or 3.2 in power. Foreground residuals are sub-dominant to the noise and create negligible bias.
@@
Gravitational lensing of the dominant E-mode polarization creates a B-mode foreground. The PIXIE 1.6 angular resolution does not permit de-lensing using only PIXIE data; any de-lensing must use external data (e.g. from CIB anisotropy or other CMB missions). Table 2 shows PIXIE B-mode limits for different levels of de-lensing, where the parameter $A_{\rm lens}$ shows the amplitude of the remaining (un-corrected) lensing signal. If no de-lensing is performed ($A_{\rm lens}=1$), the baseline PIXIE mission achieves $\sigma(r)=0.98\times 10^{-3}$ for a 95% confidence upper limit $r_{95}=2.86\times 10^{-3}$. Residual foregrounds contribute a bias $r_{\rm fg}=0.41\times 10^{-3}$, small compared to the noise. Neither the B-mode uncertainty nor the foreground residuals depend significantly on the assumed level of de-lensing. The fraction of mission observations devoted to polarization (calibrator stowed so that both beams view the sky) can be changed throughout the mission. An extended mission with a total of 7 years of polarization observation would improve the B-mode limits to $\sigma(r)=0.4\times 10^{-3}$ and $r_{95}=1.1\times 10^{-3}$, with foreground bias $r_{\rm fg}=0.08\times 10^{-3}$. With modest delensing ($A_{\rm lens}=0.5$), the 95% confidence upper limits improve to $2.6\times 10^{-3}$ for the baseline mission and $0.9\times 10^{-3}$ for an extended mission.

### v2

@@ 1 Introduction
Observations of cosmological radiation backgrounds have played a key role in our understanding of the universe. Measurements of the CMB frequency spectrum [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26] are consistent with a single blackbody over a range of 3 decades in frequency and 4 decades in intensity (Figure 1). Data from the COBE/FIRAS spectrometer determine the monopole temperature $T_{0}=2.72548\pm 0.00057$ K [27]. The FIRAS limits on deviations from a blackbody spectrum (spectral distortions) at levels $\Delta I/I<5\times 10^{-5}$ [25] support the thermal hot big bang model while ruling out alternatives such as steady state models [28]. Measurements of CMB anisotropies in temperature and polarization have provided insight into the contents of the universe and their evolution from primordial density perturbations to matter clustering, reionization, and the growth of large scale structure, consistent with a single 7-parameter cosmological model ($\Lambda$CDM, [29, 30] ).
@@
The Primordial Inflation Explorer (PIXIE) is a proposed space mission to measure the spectrum and polarization of the cosmic microwave background and astrophysical foregrounds on angular scales of 1 and larger [37, 38, 39]. Operating from the second Sun-Earth Lagrange point (L2), it would survey the entire sky in intensity and linear polarization at integrated sensitivity of a few Jy sr-1 in 300 spectral channels from 28 GHz to 6 THz (Figure 2).
PIXIE was first proposed as a NASA medium-class Explorer in 2011 [37]. Previous papers have described the PIXIE detectors [40], optics [41, 42], calibration [38], and systematic error budget [43, 39]. In this paper, we review the PIXIE instrument design, describe changes since the initial 2011 mission concept, and highlight the anticipated scientific results from its baseline 2-year mission. Section 2 describes the updated mission design, sensitivity, and foreground subtraction. Section 3 reviews the principal science goals, updates the predicted sensitivity for cosmological parameters, and provides new forecasts for measurements of the cosmic infrared background, Galactic emission lines, and the star formation rate density at redshifts $z<5$. Section 4 discusses the limitations of the single-FTS instrument configuration and outlines the sensitivity gains possible with a larger “Super-PIXIE” mission. Separate appendices detail the data sampling and interferogram apodization, outline the principal design features responsible for the sensitivity improvements relative to COBE/FIRAS, and present the full noise curves for the baseline PIXIE mission.
@@ 2 PIXIE Mission
Figure 3: Left: Exploded view of the PIXIE observatory. Nested, passively cooled shells surround a polarizing Fourier transform spectrometer, which is actively cooled to $2.725$ K to match the CMB monopole emission. Right: PIXIE’s optical design measures both intensity and polarization from microwave to THz frequencies. As the central mirror pair moves, the detectors measure a fringe pattern proportional to the Fourier transform of the difference between one linear polarization from beam A and the orthogonal polarization from beam B.
The PIXIE instrument (Figure 3) is a polarizing Fourier transform spectrometer (FTS) with two input ports and two output ports. The input ports are fed by two off-axis telescopes that produce twin beams co-pointed along the spacecraft spin axis. Each telescope has a primary mirror $550$ mm in diameter. Folding flats and secondary mirrors route the beams to the FTS while retaining orthogonal polarization alignment between the two beams. Within the FTS a set of five transfer mirror pairs, each imaging the previous mirror to the following one, shuttles the radiation through a series of polarizing wire grids. Polarizer A transmits one polarization and reflects the other, separating each beam into orthogonal polarization states Polarizer B, oriented $45^{\circ}$ relative to Polarizer A, mixes the polarization states. A Mirror Transport Mechanism (MTM) moves the central mirror pair ±$\pm 4$ mm to produce an optical phase delay of up to ±$\pm 15$ mm. The phase-delayed beams re-combine (interfere) at Polarizer C. Polarizer D (aligned to Polarizer A) splits the beams again and routes them to two multi-moded concentrators. Each concentrator has a square aperture to preserve linear polarization [41, 42] and contains a pair of orthogonal polarization-sensitive bolometers mounted back-to-back. The square aperture is aligned with respect to the plane of polarization from Polarizer D to enforce symmetry between orthogonal linear polarizations [42]. The four bolometers are cooled to 0.1 K [44] to achieve background-limited sensitivity. The optical spectrum is the Fourier conjugate of the detector’s time-ordered data across the mirror sweep. The mirror stroke length sets the optical frequency resolution while the detector sampling rate determines the highest optical frequency.
@@
The baseline 2-year mission assumes that 45% of observing time is spent with the calibrator deployed (sensitive to spectral distortions and polarization) and 45% with the calibrator stowed (sensitive to polarization only). The remaining 10% conservatively accounts for time not used for science observations (MTM turnaround at the end of each stroke, thermal settling after setpoint changes, etc). The calibrator deployment schedule can be altered during the mission to provide more or less integration time in each mode. Mission operations further assume observations at two different MTM stroke lengths, providing additional sensitivity to low-frequency foregrounds.
The PIXIE instrument and mission have evolved since the initial 2011 concept [37]. The 2011 version assumed observations from a 660 km polar sun-synchronous low-Earth orbit, while the current version will observe from the second Sun-Earth Lagrange point. The L2 orbit improves thermal stability for the observatory, reduces emission from the Earth and Moon to negligible levels, and by decoupling the great circle scan motion from the low-Earth orbital period simplifies synchronization of the interferogram mirror stroke, observatory spin, and great circle scan. Within the instrument, the optical path through the FTS has been shortened, replacing the previous corner reflector MTM with a pair of powered mirrors. The 2011 concept assumed that all observations would use the same MTM stroke, and thus the same set of frequency channels. To obtain additional low-frequency channels, the current design assumes a combination of observations with short and long MTM strokes (§$\S$2.1). The resulting sensitivity curves correct a factor-of-two error from the 2011 estimate (Eq. 2.4) and now include the frequency-dependent effects of the reflective backshort in the detector integrating cavity. The baseline mission has been shortened from 4 years to 2 years, based on the observing time required for a definitive measurement of the cosmological distortions from Compton scattering.
@@ 2.1 Sensitivity
Figure 6: Solid lines show the spectra of the principal foregrounds at the median amplitude for Galactic latitude $|b|>20^{\circ}$. Dotted and dashed lines show the cosmological signals, with negative amplitudes dashed. The black/red curve shows the PIXIE noise in each spectral channel for the baseline 2-year mission, averaged over the cleanest 70% of the sky. The rise in noise at frequencies below 40 GHz reflects the smaller fraction of observing time spent with the MTM in long-stroke mode. PIXIE’s sensitivity to cosmological signals is limited by competing emission from astrophysical foregrounds (see text).
The PIXIE channel width is determined by the need to separate cosmological signals from competing foregrounds (§$\S$2.2). Channel widths $\Delta\nu\sim 20$ GHz providing spectral resolution $\nu/\Delta\nu\sim 15$ near the CMB peak at 270 GHz do not provide sufficient resolution to fit multiple foreground components (synchrotron, free-free, and anomalous microwave emission) at frequencies below 70 GH, while channel widths $\Delta\nu<10$ GHz appropriate for low-frequency foregrounds degrade the sensitivity to CMB signals at higher frequencies. As a compromise, PIXIE spends 70% of the observing time in “short-stroke” mode using broad channel widths for continuum CMB signals, and 30% in “long-stroke” mode with narrow widths for low-frequency foregrounds. It is convenient to set all channel widths commensurate with the CO J=1-0 line at 115.3 GHz within the Galaxy, $\Delta\nu=\nu_{\rm CO}/M$ so that every $M^{\rm th}$ channel is centered on a Galactic CO line. For measurements of spectral distortions, PIXIE will use $M=6$ ($\Delta\nu=19.2$ GHz) and $M=12$ ($\Delta\nu=9.6$ GHz). The free-free and anomalous microwave foreground emission are not polarized at relevant levels, allowing use of wider channel widths $M=3$ ($\Delta\nu=38.4$ GHz) and $M=6$ ($\Delta\nu=19.2$ GHz) for polarization measurements. Appendix C lists the resulting channel sensitivities.
@@ 2.2 Foreground Subtraction
Table 1 shows the signal to noise ratio for the cosmological signals estimated using a Fisher analysis of the noise levels integrated over the cleanest 70% of the sky. We adopt fiducial values $y=1.77\times 10^{-6}$ and $kT_{e}=1.245$ keV predicted for the integrated signal from Compton scattering from groups and clusters of galaxies [52] and $\mu=2\times 10^{-8}$ for the dissipation of primordial density perturbations from a power-law distribution consistent with the scale-invariant spectrum seen in primary CMB anisotropies [53]. Results are shown for the baseline PIXIE mission (two years of observations with 45% of observations devoted to spectral distortions, yielding 7.6 months of integration for monopole distortions within the cleanest 70% of the sky) as well as an extended mission with a total of 7 years integration time in spectral distortion mode.
@@
aNot fitted
In the absence of foreground emission, the baseline PIXIE mission would detect the cosmological signals at high confidence: 3101 standard deviations for $y$, 80 standard deviations for $kT_{e}$, and 2.9 standard deviations for $\mu$. The need to identify and subtract foreground emission degrades these values. A minimal fit using just the PIXIE mission data with no foreground priors provides a robust detection of the predicted Compton signal (18 standard deviations for $y$ and 2 standard deviations for $kT_{e}$) but falls short of a detection of the chemical potential predicted in the standard $\Lambda$CDM model from Silk damping of a power-law distribution of primordial density perturbations (§$\S$3.2). We investigate the ability to improve these limits by repeating the analysis while placing priors on selected foreground parameters. PIXIE data at frequencies above $\sim$600 GHz place tight constraints on the amplitude and spectra of the thermal dust and CIB foregrounds and are unlikely to be superseded by other data. The sensitivity at lower frequencies is limited by the fixed channel width imposed by Fourier transform spectroscopy (§$\S$2.1) and can be improved by including information from measurements at lower frequencies. We repeat the Fisher analysis placing Gaussian priors of 10% or 1% on the synchrotron and free-free parameters. Placing 10% priors on the low-frequency foregrounds provides a five-fold improvement in the sensitivity to cosmological signals, but still falls short of a detection for the minimal $\mu$ distortion from a $\Lambda$CDM cosmology. Even the combination of tighter (1%) priors over a 7-year extended mission produces uncertainty larger than the predicted chemical potential distortion. Obtaining additional sensitivity to low-frequency foregrounds would require additional instrumentation to reduce photon noise at frequencies below $\sim$150 GHz, extend measurements to frequencies below 30 GHz, or both. §$\S$4 discusses possible enhancements for such a “Super-PIXIE” design.
@@
Foreground subtraction in polarization is less complex, as only the synchrotron and thermal dust foregrounds are expected to have significant polarization. Figure 7 shows PIXIE’s polarization capability using a Needlet Internal Linear Combination (NILC) component separation [54, 55, 56, 57]). We generate a noiseless full-sky foreground model using the Python Sky Model d1s1 variant [58] and add cosmological signals from E-modes, lensing B-modes, and primordial B-modes at amplitude $r=5\times 10^{-3}$. All signals are convolved with a 1.6 full width at half maximum Gaussian beam before adding noise within each PIXIE frequency channel. We complement the PIXIE frequency coverage using the Planck polarization maps, then evaluate the resulting sky maps on the cleanest 70% of the sky, using a Fisher analysis to estimate the uncertainty $\sigma(r)$ and a likelihood analysis to determine the 95% confidence upper limit $r_{95}$ [59]. As with spectral distortions, the foreground subtraction increases the effective noise with noise amplitude (after component separation) worse by a factor 1.8 in amplitude or 3.2 in power. Foreground residuals are sub-dominant to the noise and create negligible bias.
@@
Gravitational lensing of the dominant E-mode polarization creates an additional cosmological foreground for measurements of B-mode polarization from primordial tensor perturbations. The power spectrum for PIXIE B-mode noise is close to the amplitude for the un-corrected lensing foreground; roughly 70% of the weight for B-mode polarization analysis results from multipoles $\ell\leq 10$ where the primordial B-mode signal pulls away from the lensing foreground. The PIXIE 1.6 angular resolution does not permit de-lensing using only PIXIE data; any de-lensing must use external data (e.g. from CIB anisotropy or other CMB missions). Table 2 shows PIXIE B-mode limits for different levels of de-lensing, where the parameter $A_{\rm lens}$ shows the amplitude of the remaining (un-corrected) lensing signal. If no de-lensing is performed ($A_{\rm lens}=1$), the baseline PIXIE mission achieves $\sigma(r)=0.98\times 10^{-3}$ for a 95% confidence upper limit $r_{95}=2.86\times 10^{-3}$. Residual foregrounds contribute a bias $r_{\rm fg}=0.41\times 10^{-3}$, small compared to the noise. Neither the B-mode uncertainty nor the foreground residuals depend significantly on the assumed level of de-lensing. The fraction of mission observations devoted to polarization (calibrator stowed so that both beams view the sky) can be changed throughout the mission. An extended mission with a total of 7 years of polarization observation would improve the B-mode limits to $\sigma(r)=0.4\times 10^{-3}$ and $r_{95}=1.1\times 10^{-3}$, with foreground bias $r_{\rm fg}=0.08\times 10^{-3}$. With modest delensing ($A_{\rm lens}=0.5$), the 95% confidence upper limits improve to $2.6\times 10^{-3}$ for the baseline mission and $0.9\times 10^{-3}$ for an extended mission.
@@
The above forecasts assume a simple foreground model (d1s1) from widely used PySM sky model. Recent Planck data suggest that variation of the thermal dust spectral across the sky may be larger than assumed in this model. To investigate the impact of additional complexity in the foreground model, we have run the NILC code on more recent models implemented in PySM3, which use template maps and statistics of foreground spectra from Planck public data release 3 [57, 60]. The medium complexity model (PySM3 d10s5) is similar to the baseline low-complexity model except that the dust model d10 shows significantly more spatial variability of the dust spectrum in mid-to-high Galactic latitudes. The high-complexity sky model (PySM3 d12s7a2) adds 2% fractional polarization for AME, spectral curvature for synchrotron emission, and adopts a 6-layer multi-component model for polarized dust emission [61]. The medium and high complexity models are expected to be valid only within the limited frequency range 1–1000 GHz. We thus restrict the NILC analysis to PIXIE channels below 1000 GHz, but include an external channel with beam width and sensitivity comparable to WMAP K band as an additional tracer of polarized synchrotron.
Over the cleanest 70% of the sky, the medium or high complexity models degrade the sensitivity $\sigma(r)$ by a factor 2–3 compared to the baseline model. The reduced sensitivity is driven almost entirely by complexity in the dust model. Although use of the PySM3 models restricts analysis to frequencies below 1000 GHz, PIXIE will produce 52 frequency channels sensitive to polarized dust at frequencies 1–3.5 THz, with median signal-to-noise ratios 25–600 in individual high-latitude voxels within this frequency range. Observations with the calibrator deployed provide additional, independent data for both polarized and unpolarized foregrounds. A full treatment of foreground modeling with PIXIE is beyond the scope of this paper and is deferred to a future work.
3 Science Goals
3.1 Inflation